OpenStudy (anonymous):

use the trigonometric function vngles given in this section to evaluate each experssixpression such as cot"2 90° means (cot90°)"2 , which is equal to 0"2 = 0. sec"2 180° -35in"2 360° +cos180°

OpenStudy (freckles):

what? I don't understand... do you mean \[\sec^2(180^o)-3 \sin^2(360^o)+\cos(180^o)\]?

OpenStudy (anonymous):

Yes :)

OpenStudy (freckles):

do you know how to evaluate cos(180)?

OpenStudy (anonymous):

yes it is -1... according of what my teacher said..

OpenStudy (freckles):

\[\sec^2(180)-3 \sin^2(360)+\cos(180) \\ \frac{1}{(\cos(180))^2}-3(\sin(360))^2+\cos(180)\]

OpenStudy (freckles):

all you really need to do for this question is know how to evaluate cos(180) and sin(360) the rest his order of operations

OpenStudy (anonymous):

but why do you convert sec to cos? and what do I do with the 2's?

OpenStudy (freckles):

sec(x)=1/cos(x)

OpenStudy (freckles):

cos and sec are reciprocal functions of one another

OpenStudy (freckles):

can you tell me what cos(180) and sin(360) equals?

OpenStudy (anonymous):

cos= -1 and sin =0

OpenStudy (freckles):

so we are just going to replace those items real quick like this: \[\sec^2(180)-3 \sin^2(360)+\cos(180) \\ \frac{1}{(\cos(180))^2}-3(\sin(360))^2+\cos(180) \\ \frac{1}{(-1)^2}-3(0)^2+-1 \]

OpenStudy (freckles):

now just apply order of operations

OpenStudy (anonymous):

all equal -3?

OpenStudy (freckles):

1-3(0)-1 isn't -3

OpenStudy (anonymous):

0?

OpenStudy (freckles):

seems great me so if cos(180)=-1 then squaring both sides gives you (cos(180))^2=(-1)^2=1 and if sin(360)=0 then squaring both sides gives you (sin(360))^2=(0)^2=0 so yes we have 1-3(0)-1=1-0-1=1-1=0

OpenStudy (anonymous):

Ya ! Thank you so much for heping me @freckles have a nice day ;)